Suppose Jada Corp’s stock beta is 1.4. If the current market risk premium is 8% and the risk-free rate is 2%, what is Jada Corp’s required return?

10.40%

11.20%

14.00%

13.20%

None of these

The expected return of the S&P 500 Index is 10%. According to stock analysts, Ingrid Corp stock has an expected return of 15%. You should buy this stock if you believe the following:

You should buy the stock, regardless of the stock’s beta.

You should never buy the stock.

Ingrid Corp’s beta is 2.1.

Not enough information given to make an informed decision.

Ingrid Corp’s beta is 1.6.
Answer #1

1. Option (d) is correct

As per CAPM equation, required return of the stock is given by:

Required return = Risk free rate + Beta * Market risk premium

Given: Risk free rate = 2%, Market risk premium = 8%, Beta = 1.4

Putting the given values in the above equation, we get,

Required rate of return = 2 + (1.4 * 8)

Required rate of return = 2 + 11.2 = 13.2

2. Option (a) is correct

As expected return of the stock (15%) is higher than the market return (10%), we should buy the stock. Expected return of the stock reflects its beta and market risk premium.